# Euler characteristics of the universal Picard stack

### v1

@@ Abstract.
We study $\mathbb{S}_{n}$-equivariant topological and weight-graded compactly-supported Euler characteristics of the universal Picard stack $\mathrm{Pic}_{g,n}^{d}\to\mathcal{M}_{g,n}$ of degree-$d$ line bundles over $\mathcal{M}_{g,n}$. We prove that in the weight-zero and topological cases, the generating function for Euler characteristics of $\mathrm{Pic}_{g,n}^{d}$ is obtained from the corresponding one for $\mathcal{M}_{g,n}$ by an extremely simple combinatorial transformation. This lets us deduce closed formulas for the two generating functions, taking as input the Chan–Faber–Galatius–Payne formula in the weight-zero case and Gorsky’s formula in the topological case. As a corollary, we also obtain a closed formula for the topological Euler characteristic of $\mathrm{Pic}^{d}_{g}$. Our weight-zero calculation is a corollary of a general result passing from the weight-graded Euler characteristics of $\mathcal{M}_{g,n}$ to those of $\mathrm{Pic}_{g,n}^{d}$.
@@ 1. Introduction
Fix integers $g\geq 2$ and $n,d\geq 0$, and consider the universal Picard stack
$\mathrm{Pic}^{d}_{g,n}\to\mathcal{M}_{g,n},$
@@
be the $k$th graded piece of the compactly-supported cohomology. We are interested in the $\mathbb{S}_{n}$-equivariant weight-$k$ compactly-supported Euler characteristic
@@
is the ring of degree-completed symmetric functions, and $\operatorname{ch}_{n}(V)$ denotes the Frobenius characteristic of a finite-dimensional $\mathbb{S}_{n}$-representation $V$. The weight-$k$ Euler characteristic of $\mathrm{Pic}^{d}_{g,n}$ does not depend on $d$ [19, Lemma 4.3], so we will write
@@
$\mathsf{J}_{g}^{k}:=\sum_{n\geq 0}\chi_{k}^{\mathbb{S}_{n}}(\mathrm{Pic}_{g,n})\quad\mbox{and}\quad\mathsf{Z}_{g}^{k}:=\sum_{n\geq 0}\chi_{k}^{\mathbb{S}_{n}}(\mathcal{M}_{g,n}).$
Our first main theorem passes from $\mathsf{Z}_{g}^{k}$ to $\mathsf{J}_{g}^{k}$. It implies a closed formula for $\mathsf{J}_{g}^{0}$ (Corollary 2), and a similar technique yields a closed formula for topological Euler characteristics (Theorem 3). Sample calculations are included at the end of the paper in Tables 1 and 2.
@@
The generating function $\mathsf{J}_{g}^{0}$ is obtained from the expression for $\mathsf{Z}_{g}^{0}$ in [5, Theorem 1.1] by restricting the sum to those Laurent monomials in the $P_{j}$ whose exponents sum to $1$, and then making the substitution $P_{j}\mapsto P_{j}/j$ for all $j$. Specifically,
@@
where the sum is over integers $k,m,s>0$ and $s$-tuples of positive integers $a=(a_{1},\ldots,a_{s})$ and $d=(d_{1},\ldots,d_{s})$, such that
$0<d_{1}<\cdots<d_{s}<m\quad\mbox{and}\quad d_{i}\mid m;$
$\gcd(d_{1},\ldots,d_{s})=1;$
$a_{1}+\cdots+a_{s}=k+1;$
$a_{1}d_{1}+\cdots+a_{s}d_{s}+g-1=km.$
It is rather non-obvious to me a priori that such a simple transformation should calculate $\mathsf{J}_{g}^{0}$. If we say that $P_{i}$ has degree $i$, Corollary 2 implies that $\mathsf{J}_{g}^{0}$ is a homogeneous Laurent polynomial in the $P_{i}$’s of degree 1g1-g, as is true for $\mathsf{Z}_{g}^{0}$. See Table 1 for some values of $\mathsf{J}_{g}^{0}$.
@@
$\mathsf{J}_{g}^{top}:=\sum_{n\geq 0}\chi^{\mathbb{S}_{n}}_{top}(\mathrm{Pic}_{g,n})\quad\mbox{and}\quad\mathsf{Z}_{g}^{top}:=\sum_{n\geq 2-2g}\chi^{\mathbb{S}_{n}}_{top}(\mathcal{M}_{g,n}).$
A beautiful formula for $\mathsf{Z}_{g}^{top}$ was obtained by E. Gorsky [14, Theorem 3.8]. In his notation, $1+p_{j}t^{j}$ corresponds to $P_{j}$. We find another elementary combinatorial transformation of $\mathsf{Z}_{g}^{top}$ gives $\mathsf{J}_{g}^{top}$. In this case we focus on monomials of total exponent $2$.
@@
The generating function $\mathsf{J}_{g}^{top}$ is obtained from Gorsky’s formula [14, Theorem 3.8] for $\mathsf{Z}_{g}^{top}$ by restricting the sum to Laurent monomials in the $P_{j}$ whose exponents sum to $2$, and then making the substitution $P_{j}\mapsto P_{j}/j$ for all $j$. Explicitly,
@@
Just like $\mathsf{Z}_{g}^{top}$, the expression for $\mathsf{J}_{g}^{top}$ is a Laurent polynomial in the $P_{i}$’s which is homogeneous of degree 22g2-2g. As an immediate corollary, we get a closed formula for the topological Euler characteristic $\chi_{top}(\mathrm{Pic}_{g})$.
@@
We have
$\chi_{top}\left(\mathrm{Pic}_{g}\right)=\sum_{r=2}^{4g+2}\sum_{s=3}^{2g+2}\sum_{k_{1},\ldots,k_{r}}\left((-1)^{s-3}(s-3)!\frac{N(r;k_{1},\ldots,k_{r-1})}{\prod_{i=1}^{r-1}i^{k_{i}}k_{i}!}\cdot r^{s-3}\right),$
where the sum is over the same $k_{i}$ as in Theorem 3.
See Table 2 for some values of $\mathsf{J}_{g}^{top}$. The number $\chi(\mathrm{Pic}_{g}^{0})$ is obtained by setting all of the $P_{i}$ to $1$.
@@ 1.1. Remarks on the formulas
Note that for $n>2g+2$, the map $\mathrm{Pic}_{g,n}^{d}\to\mathcal{M}_{g,n}$ is an abelian fibration of varieties, which means $\chi_{top}(\mathrm{Pic}_{g,n}^{d})=0$ in this range. This property is manifest in Theorem 3, since the generating function $\sum_{n}\chi_{top}(\mathrm{Pic}_{g,n})t^{n}/n!$ for numerical Euler characteristics is obtained by sending $P_{1}\to(1+t)$ and $P_{i}\to 1$ for $i>0$. One verifies that there are no negative powers of $P_{1}$ appearing in the formula and that the maximum power of $P_{1}$ which appears is $2g+2$, which then implies the vanishing statement. However, this vanishing does not hold $\mathbb{S}_{n}$-equivariantly: the invariant $\chi_{top}^{\mathbb{S}_{n}}$ is still interesting even for $n>2g+2$.
More surprising to me is that an analogous property holds for the numerical weight-zero Euler characteristic $\chi_{0}(\mathrm{Pic}_{g,n})$. Via a similar analysis of the monomials appearing in the formula, we find that it vanishes for $n>g+1$, but not $\mathbb{S}_{n}$-equivariantly. It would be interesting to find a explanation for this property which comes from tropical geometry; I expect that it is related to tropical Jacobians of metric graphs [1].
@@ 1.2. Techniques
To calculate with relative symmetric powers, we will start with a generating function for Euler characteristics of $\mathcal{M}_{g,n}$, and then make a series of transformations using symmetric function theory. Each step has an interpretation in terms of allowing or forbidding groups of marked points to coincide, or in terms of taking symmetric group quotients. When these transformations are composed, we are led to the algebra map $\mathcal{T}$ of (1.1). The geometry behind these transformations was studied in [17], in the context of certain moduli spaces of weighted pointed stable curves [16].
@@ 1.3. Related work
In [19], T. Song and I studied the weight-graded Euler characteristics characteristics of $\mathrm{Pic}_{g,n}$, towards understanding those of $\mathcal{M}_{g,n}(\mathbb{P}^{r},d)$. In particular, Proposition E in loc. cit. determines the generating function for $\mathbb{S}_{n}$-equivariant Hodge–Deligne polynomials of $\mathrm{Pic}_{g,n}$ in terms of the corresponding generating function for $\mathcal{M}_{g,n}$. It is not clear how to extract the simple formulas in the present work directly from [19, Proposition E], but I emphasize that the geometric insight used here is the same: it is the algebraic approach, particularly the definition and analysis of the operator $\mathcal{T}$ of (1.1), that is different.
@@ 2. Transforms of symmetric functions
Let $\mathrm{Cl}_{\mathbb{Q}}(\mathbb{S}_{n})$ denote the vector space space of $\mathbb{Q}$-valued class functions on $\mathbb{S}_{n}$. There is a natural identification
@@ 2.1. Bisymmetric functions
where $\otimes$ means the completed tensor product, and we have identified $p_{i}=p_{i}\otimes 1$ and $q_{i}=1\otimes q_{i}$.
@@ 3. Proofs of the main theorems
Let $\mathcal{C}_{g}^{n}\to\mathcal{M}_{g}$ be the $n$th fibered power of the universal curve over $\mathcal{M}_{g}$. The stack $\mathcal{C}_{g}^{n}$ parameterizes smooth curves of genus $g$ with $n$ ordered, not-necessarily-distinct marked points, and is naturally identified with the stack $\mathcal{M}_{g,\varepsilon^{n}}$ of smooth curves in Hassett’s compactification $\overline{\mathcal{M}}_{g,\varepsilon^{n}}$ of $\mathcal{M}_{g,n}$, for any $\varepsilon\in\mathbb{Q}_{>0}$ with $\varepsilon<1/n$. We thus obtain the following lemma from [17, Theorem A].
@@
$\mathsf{Z}_{g}^{k}\circ\operatorname{Exp}(p_{1})=\sum_{n\geq 0}\chi_{k}^{\mathbb{S}_{n}}(\mathcal{C}_{g}^{n}).$
@@
(3.2) $\sum_{m,n\geq 0}\chi_{k}^{\mathbb{S}_{m}\times\mathbb{S}_{n}}(\mathcal{M}_{g,m|n})=\Delta(\mathsf{Z}_{g}^{k})\circ_{2}\operatorname{Exp}(q_{1}),$
where we have again set
@@
$\sum_{n\geq 0}\chi_{k}^{\mathbb{S}_{n}}(\mathcal{C}_{g}^{n})=\mathsf{Z}_{g}^{k}\circ\operatorname{Exp}(p_{1}),$
@@
From (2.8) we obtain the following corollary.
@@
$\mathrm{inv}\Delta(\mathsf{Z}_{g}^{k}\circ\operatorname{Exp}(p_{1}))=\sum_{m,n\geq 0}\chi_{k}^{\mathbb{S}_{n}}([\mathcal{C}_{g}^{m+n}/\mathbb{S}_{m}])x^{m}.$
Recall that the projective bundle $\mathcal{S}_{g,n}^{m}$ of (1.2) is defined for $m\geq 2g-1$, and its rational cohomology coincides with that of the stack quotient $\mathcal{C}_{g,n}^{m}/\mathbb{S}_{m}]$. For ease of exposition, we formally set
@@
for $0\leq m\leq 2g-2$. The following lemma follows from the same reasoning as [17, Proposition 4.3], or [12, Theorem 3.2]: the point is that the operation $\circ\operatorname{Exp}(p_{1})$ allows marked points to coincide.
@@
$\left(\sum_{m,n\geq 0}\chi_{k}^{\mathbb{S}_{n}}(\mathcal{S}_{g,n}^{m})x^{m}\right)\circ\operatorname{Exp}(p_{1})=\sum_{n\geq 0}\chi_{k}^{\mathbb{S}_{n}}([\mathcal{C}_{g}^{m+n}/\mathbb{S}_{m}])x^{m}.$
Noting that $\operatorname{Log}(1+p_{1})$ is the inverse of $\operatorname{Exp}(p_{1})$ under plethysm, we obtain the following corollary by combining Corollary 3.3 and Lemma 3.4.
@@
$\sum_{n\geq 0}\chi^{\mathbb{S}_{n}}_{k}(\mathcal{S}_{g,n}^{m})x^{m}=\mathrm{inv}\Delta(\mathsf{Z}_{g}^{k}\circ\operatorname{Exp}(p_{1}))\circ\operatorname{Log}(1+p_{1}).$

### v2

@@ Abstract.
We study $\mathbb{S}_{n}$-equivariant weight-graded and topological Euler characteristics of the universal Picard stack $\mathrm{Pic}_{g,n}^{d}\to\mathcal{M}_{g,n}$ of degree-$d$ line bundles over $\mathcal{M}_{g,n}$. We prove that in the weight-zero and topological cases, the generating function for Euler characteristics of $\mathrm{Pic}_{g,n}^{d}$ is obtained from the corresponding one for $\mathcal{M}_{g,n}$ by an extremely simple combinatorial transformation. This lets us deduce closed formulas for the two generating functions, taking as input the Chan–Faber–Galatius–Payne formula in the weight-zero case and Gorsky’s formula in the topological case. As an immediate corollary, we obtain closed formulas for the weight-zero and topological Euler characteristics of $\mathrm{Pic}^{d}_{g}$. Our weight-zero calculations follow from a general result passing from the weight-graded Euler characteristics of $\mathcal{M}_{g,n}$ to those of $\mathrm{Pic}_{g,n}^{d}$.
@@ 1. Introduction
Fix integers $g\geq 1$ and $n,d\geq 0$ such that $(g,n)\neq(1,0)$. Consider the universal Picard stack
$\mathrm{Pic}^{d}_{g,n}\to\mathcal{M}_{g,n}$
@@
be the $k$th weight-graded piece of the compactly-supported cohomology. We are interested in the $\mathbb{S}_{n}$-equivariant weight-$k$ compactly-supported Euler characteristic
@@
is the ring of degree-completed symmetric functions and $\operatorname{ch}_{n}(V)$ denotes the Frobenius characteristic of a finite-dimensional $\mathbb{S}_{n}$-representation $V$. The weight-$k$ Euler characteristic of $\mathrm{Pic}^{d}_{g,n}$ does not depend on $d$ [19, Lemma 4.3], so we will write
@@
(1.1) $\mathsf{J}_{g}^{k}:=\sum_{n}\chi_{k}^{\mathbb{S}_{n}}(\mathrm{Pic}_{g,n})\quad\mbox{and}\quad\mathsf{Z}_{g}^{k}:=\sum_{n}\chi_{k}^{\mathbb{S}_{n}}(\mathcal{M}_{g,n});$
the sums are taken over $n\geq 1$ if $g=1$ and over $n\geq 0$ otherwise. Our main result Theorem 1 passes from $\mathsf{Z}_{g}^{k}$ to $\mathsf{J}_{g}^{k}$. It implies a closed formula for $\mathsf{J}_{g}^{0}$ (Corollary 2), and a similar technique yields a closed formula for topological Euler characteristics (Theorem 3). Sample calculations are included at the end of the paper in Tables 1 and 2.
@@
Suppose $g\geq 2$. The generating function $\mathsf{J}_{g}^{0}$ is obtained from the expression for $\mathsf{Z}_{g}^{0}$ in [5, Theorem 1.1] by restricting the sum to those Laurent monomials in the $P_{j}$ whose exponents sum to $1$, and then making the substitution $P_{j}\mapsto P_{j}/j$ for all $j$. Specifically,
@@
where the sum is over integers $k,m,s>0$ and $s$-tuples of positive integers $a=(a_{1},\ldots,a_{s})$ and $d=(d_{1},\ldots,d_{s})$ with $\gcd(d_{1},\ldots,d_{s})=1$ and $d_{i}\mid m$ for each $i$, such that $0<d_{1}<\cdots<d_{s}<m$, $\sum_{i}a_{i}=k+1$, and $\sum_{i}a_{i}d_{i}=km+1-g$.
It is rather non-obvious to me a priori that such a simple transformation should calculate $\mathsf{J}_{g}^{0}$. If we say that $P_{j}$ has degree $j$, Corollary 2 implies that $\mathsf{J}_{g}^{0}$ is a homogeneous Laurent polynomial in the $P_{j}$’s of degree 1g1-g, as is true for $\mathsf{Z}_{g}^{0}$. See Table 1 for some values of $\mathsf{J}_{g}^{0}$.
@@
$\mathsf{J}_{g}^{top}:=\sum_{n}\chi^{\mathbb{S}_{n}}_{top}(\mathrm{Pic}_{g,n})\quad\mbox{and}\quad\mathsf{Z}_{g}^{top}:=\sum_{n}\chi^{\mathbb{S}_{n}}_{top}(\mathcal{M}_{g,n});$
the sums are over the same $n$ as in (1.1). When $g\geq 2$, a beautiful formula for $\mathsf{Z}_{g}^{top}$ was obtained by E. Gorsky [14, Theorem 3.8]. In his notation, $1+p_{j}t^{j}$ corresponds to $P_{j}$. We find that another elementary combinatorial transformation of $\mathsf{Z}_{g}^{top}$ gives $\mathsf{J}_{g}^{top}$. In this case we focus on monomials of total exponent $2$.
@@
Suppose $g\geq 2$. The generating function $\mathsf{J}_{g}^{top}$ is obtained from Gorsky’s formula [14, Theorem 3.8] for $\mathsf{Z}_{g}^{top}$ by restricting the sum to Laurent monomials in the $P_{j}$ whose exponents sum to $2$, and then making the substitution $P_{j}\mapsto P_{j}/j$ for all $j$. Explicitly,
@@
Just like $\mathsf{Z}_{g}^{top}$, the expression for $\mathsf{J}_{g}^{top}$ is a Laurent polynomial in the $P_{j}$’s which is homogeneous of degree 22g2-2g. See Table 2 for some values of $\mathsf{J}_{g}^{top}$. Analogous formulas for $\mathsf{J}_{1}^{0}$ and $\mathsf{J}_{1}^{top}$ can also be obtained, though we do not state them explicitly here. See [5, Proposition 1.5] for the weight-zero case, and note that [10, Equation (5.5)] may be specialized to the topological case.
From the definitions of the relevant generating functions, the preceding results give formulas for the weight-zero and topological Euler characteristics of $\mathrm{Pic}_{g}$.
@@
For $g\geq 2$, closed formulas for $\chi_{0}(\mathrm{Pic}_{g})$ and $\chi_{top}(\mathrm{Pic}_{g})$ are obtained from $\mathsf{J}_{g}^{0}$ and $\mathsf{J}_{g}^{top}$ respectively by the substitution $P_{j}\to 1$ for all $j$.
@@ 1.1. Remarks on the formulas
Note that for $n>2g+2$, the map $\mathrm{Pic}_{g,n}^{d}\to\mathcal{M}_{g,n}$ is an abelian fibration of varieties, which means $\chi_{top}(\mathrm{Pic}_{g,n}^{d})=0$ in this range. This property is manifest in Theorem 3, since the generating function $\sum_{n}\chi_{top}(\mathrm{Pic}_{g,n})t^{n}/n!$ for numerical Euler characteristics is obtained by sending $P_{1}\to(1+t)$ and $P_{j}\to 1$ for $j>0$. One verifies that there are no negative powers of $P_{1}$ appearing in the formula and that the maximum power of $P_{1}$ which appears is $2g+2$, which then implies the vanishing statement. However, this vanishing does not hold $\mathbb{S}_{n}$-equivariantly: the invariant $\chi_{top}^{\mathbb{S}_{n}}$ is still interesting even for $n>2g+2$.
More surprising to me is that an analogous property holds for the numerical weight-zero Euler characteristic $\chi_{0}(\mathrm{Pic}_{g,n})$. Via a similar analysis of the monomials appearing in the formula, we find that it vanishes for $n>g+1$, but not $\mathbb{S}_{n}$-equivariantly. It would be interesting to find an explanation for this property which comes from tropical geometry; I expect that it is related to tropical Jacobians of metric graphs [1].
@@ 1.2. Techniques
To calculate with relative symmetric powers, we will start with a generating function for Euler characteristics of $\mathcal{M}_{g,n}$, and then make a series of transformations using symmetric function theory. Each step has an interpretation in terms of allowing or forbidding groups of marked points to coincide, or in terms of taking symmetric group quotients. When these transformations are composed, we are led to the algebra map $\mathcal{T}$ of (1.2). The geometry behind these transformations was studied in [17], in the context of Hassett’s moduli spaces of weighted pointed stable curves [16].
@@ 1.3. Related work
In [19], T. Song and I studied the weight-graded Euler characteristics characteristics of $\mathrm{Pic}_{g,n}$, towards understanding those of $\mathcal{M}_{g,n}(\mathbb{P}^{r},d)$. In particular, Proposition E in loc. cit. determines the generating function for $\mathbb{S}_{n}$-equivariant Hodge–Deligne polynomials of $\mathrm{Pic}_{g,n}$ in terms of the corresponding generating function for $\mathcal{M}_{g,n}$. It is not clear how to extract the simple formulas in the present work directly from [19, Proposition E], and the calculation herein is entirely self-contained. The geometric insight used here is the same as [19]: it is the algebraic approach, particularly the definition and analysis of the operator $\mathcal{T}$ of (1.2), that is different.
@@ 2. Transforms of symmetric functions
Let $\mathrm{Cl}_{\mathbb{Q}}(\mathbb{S}_{n})$ denote the vector space of $\mathbb{Q}$-valued class functions on $\mathbb{S}_{n}$. There is a natural identification
@@ 2.1. Bisymmetric functions
where $\otimes$ means the completed tensor product, and we have identified $p_{i}=p_{i}\otimes 1$ and $q_{i}=1\otimes p_{i}$.
@@ 3. Proofs of the main theorems
For $g\geq 2$ and $n>0$ define $\mathcal{C}_{g}^{n}\to\mathcal{M}_{g}$ be the $n$th fibered power of the universal curve over $\mathcal{M}_{g}$, and formally set $\mathcal{C}_{g}^{0}=\mathcal{M}_{g}$. The stack $\mathcal{C}_{g}^{n}$ parameterizes smooth curves of genus $g$ with $n$ ordered, not-necessarily-distinct marked points, and is naturally identified with the stack $\mathcal{M}_{g,\varepsilon^{n}}$ of smooth curves in Hassett’s compactification $\overline{\mathcal{M}}_{g,\varepsilon^{n}}$ of $\mathcal{M}_{g,n}$, for any $\varepsilon\in\mathbb{Q}_{>0}$ with $\varepsilon<1/n$; this defines $\mathcal{C}_{g}^{n}$ for $g=1$ and $n>0$ as well. We thus obtain the following lemma from [17, Theorem A].
@@
$\mathsf{Z}_{g}^{k}\circ\operatorname{Exp}(p_{1})=\sum_{n}\chi_{k}^{\mathbb{S}_{n}}(\mathcal{C}_{g}^{n}).$
where the sum is over those $n$ as in (1.1).
@@
(3.2) $\sum_{m,n}\chi_{k}^{\mathbb{S}_{m}\times\mathbb{S}_{n}}(\mathcal{M}_{g,m|n})=\Delta(\mathsf{Z}_{g}^{k})\circ_{2}\operatorname{Exp}(q_{1}),$
where the sum on the left-hand-side is over all $m,n\geq 0$ and must satisfy the additional constraint that $m+n>0$ if $g=1$. We have again set
@@
$\sum_{n}\chi_{k}^{\mathbb{S}_{n}}(\mathcal{C}_{g}^{n})=\mathsf{Z}_{g}^{k}\circ\operatorname{Exp}(p_{1}),$
@@
From (2.8) we obtain the following corollary. All sums over $m,n$ in this section are taken with the same convention as in (3.2).
@@
$\mathrm{inv}\Delta(\mathsf{Z}_{g}^{k}\circ\operatorname{Exp}(p_{1}))=\sum_{m,n}\chi_{k}^{\mathbb{S}_{n}}([\mathcal{C}_{g}^{m+n}/\mathbb{S}_{m}])x^{m}.$
Recall that the projective bundle $\mathcal{S}_{g,n}^{m}$ of (1.3) is defined for $m\geq 2g-1$, and its rational cohomology coincides with that of the stack quotient $\mathcal{C}_{g,n}^{m}/\mathbb{S}_{m}]$. We formally set
@@
for $0\leq m\leq 2g-2$. The next lemma follows from the same reasoning as [17, Proposition 4.3], or [12, Theorem 3.2]: the operation $\circ\operatorname{Exp}(p_{1})$ allows marked points to coincide.
@@
$\left(\sum_{m,n}\chi_{k}^{\mathbb{S}_{n}}(\mathcal{S}_{g,n}^{m})x^{m}\right)\circ\operatorname{Exp}(p_{1})=\sum_{m,n}\chi_{k}^{\mathbb{S}_{n}}([\mathcal{C}_{g}^{m+n}/\mathbb{S}_{m}])x^{m}.$
Since $\operatorname{Log}(1+p_{1})$ is the inverse of $\operatorname{Exp}(p_{1})$ under plethysm, we obtain the following result by combining Corollary 3.3 and Lemma 3.4.
@@
$\sum_{m,n}\chi^{\mathbb{S}_{n}}_{k}(\mathcal{S}_{g,n}^{m})x^{m}=\mathrm{inv}\Delta(\mathsf{Z}_{g}^{k}\circ\operatorname{Exp}(p_{1}))\circ\operatorname{Log}(1+p_{1}).$